OpenStudy (anonymous):

Give exact and approximate solutions to three decimal places. x^2 +8x +16=1 What are the exact solutions? x=

OpenStudy (aroub):

x=-5 , x= -3

OpenStudy (radar):

\[x ^{2}+8x+16-1=0\]\[x ^{2}-8x+15=0\](x-5)(x-3)=0 x=5.000 x=3.000

OpenStudy (radar):

To three decimal places or x=3 and x=5 exactly.

OpenStudy (anonymous):

so I'm a little confused is it plus or minus?